# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

### v3

@@ F.1 Algorithmic description
7:       $\mathcal{I}_{i}\leftarrow\{\,n\mid g_{n}=G_{j}\}$// all samples in group $G_{j}$
@@
10:       $\Omega\;\leftarrow\;\min_{i}\{\mathrm{ACC}(G_{j})\}\;\geq\;\alpha\;\wedge\;\min_{i}\{\mathrm{F1}(G_{j})\}\;\geq\;\beta$ // boolean output
@@
14:         $\Psi\;\leftarrow\;\Bigl{[}\,\max_{k}\Delta_{k}\;\leq\;\delta_{\mathrm{fair}}\Bigr{]}$ // boolean output
15:       for each group $G_{j}$ do
16:         $L_{i}(\theta)=-\mathrm{Acc}_{j}(\theta)+\kappa\,[\,\beta-\mathrm{F1}_{j}(\theta)\,]_{+}$
17:
@@
19:         Clip $\theta(G_{j})$ to stay within $[a,b]$
20:    $\Delta_{\theta}\;\leftarrow\;\max_{G_{j}}\bigl{|}\theta(G_{j})-\text{previousThresholds}[G_{j}]\bigr{|}$
21:    if ($\Delta_{\theta}<\epsilon_{\mathrm{tol}}\;\wedge\;\Omega\;\wedge\;\Psi$then
22:       break
@@
24:    $\text{iteration}\leftarrow\text{iteration}+1$
25:  return $\{\theta(G_{1}),\dots,\theta(G_{j})\}$
@@
F.2 Early stopping method for the tradeoff curve
An optimization algorithm was implemented using a function from the scipy.optimize library, aiming to determine probability thresholds that satisfy specific fairness constraints (acceptable_disparity = [1.00, 0.30, 0.25, 0.21. 0.20, 0.19, 0.15, 0.10]) while ensuring minimum performance thresholds for both performance criteria (ACC and F1). The optimization in the minimize function was performed using the L-BFGS-B method. Early stopping is facilitated by the default values of ftol and gtol in L-BFGS-B. Additionally, the learning rate was set to $10^{-3}$ to enable controlled and stable updates during optimization, and a penalty parameter was set to $10$ to strengthen the L2 penalty in the loss function. Furthermore, minimum performance thresholds for accuracy and F1-score were enforced based on the performance observed for each group when applying a static threshold of 0.5 and a ROC-based threshold. As stricter fairness constraints are imposed, these minimum thresholds are gradually relaxed.
F.3 Hyperparameter settings for the FairOPT
The application of the FairOPT is configured with a set of hyperparameters designed to ensure a balance between fairness and performance during the threshold optimization process. The learning rate is set to $10^{-3}$ to enable controlled and stable updates during optimization. The maximum number of iterations is specified as $10^{5}$ to provide sufficient time for the optimization algorithm to converge. An acceptable disparity of 0.2 is defined to regulate subgroup fairness, ensuring that performance differences across subgroups remain within an acceptable range through the relaxed fairness. Minimum thresholds for accuracy (0.25) and F1-score (0.25) are enforced to maintain baseline model performance and stability, particularly in cases of imbalanced datasets. A tolerance level of $10^{-2}$ is introduced to impose convergence criteria, ensuring the optimization process halts only when the changes in the objective function are negligible. Furthermore, a penalty parameter of 20 is applied to enforce performance constraints by penalizing significant performance disparities across subgroups. For fairness criteria, Demographic Parity (DP) was employed to test whether the generated thresholds meets the expected fairness. Also, this experiment used early stopping method for effective processing. The early stopping check through the change of the ACC, F1, and maximum DP discrepancy of each group, and terminates if all of them shows minimal change based on predefined patience.

### v4

@@ H.1 Algorithmic description
7:        $\mathcal{I}_{i}\leftarrow\{\,n\mid g_{n}=G_{j}\}$  all samples in group $G_{j}$
@@
10:        $\Omega\;\leftarrow\;\min_{i}\{\mathrm{ACC}(G_{j})\}\;\geq\;\alpha\;\wedge\;\min_{i}\{\mathrm{F1}(G_{j})\}\;\geq\;\beta$ boolean output
@@
14:           $\Psi\;\leftarrow\;\Bigl{[}\,\max_{k}\Delta_{k}\;\leq\;\delta_{\mathrm{fair}}\Bigr{]}$ boolean output
15:        end for
16:        for each group $G_{j}$ do
17:           $L_{i}(\theta)=-\mathrm{Acc}_{j}(\theta)+\kappa\,[\,\beta-\mathrm{F1}_{j}(\theta)\,]_{+}$
18:
@@
20:           Clip $\theta(G_{j})$ to stay within $[a,b]$
21:        end for
22:     end for
23:     $\Delta_{\theta}\;\leftarrow\;\max_{G_{j}}\bigl{|}\theta(G_{j})-\text{previousThresholds}[G_{j}]\bigr{|}$
24:     if ($\Delta_{\theta}<\epsilon_{\mathrm{tol}}\;\wedge\;\Omega\;\wedge\;\Psi$then
25:        break
26:     end if
@@
28:     $\text{iteration}\leftarrow\text{iteration}+1$
29:  end while
30:  return $\{\theta(G_{1}),\dots,\theta(G_{j})\}$
@@
H.2 Early stopping method for the tradeoff curve
The FairOPT algorithm was implemented using functions from the scipy.optimize library, aiming to determine probability thresholds that satisfy specific fairness constraints (acceptable_disparity = [1.00, 0.30, 0.25, 0.21, 0.20, 0.19, 0.15, 0.15]) while ensuring minimum performance thresholds for both performance criteria (ACC and F1). The optimization in the minimize function was performed using the L-BFGS-B method. The figure is based on the RoBERTa-base model, and it uses different hyperparameters to get eight results with different relaxation methods. Early stopping is facilitated by the default values of ftol and gtol in L-BFGS-B. Additionally, the learning rate was set to $10^{-3}$ to enable controlled and stable updates during optimization, and a penalty parameter was set to $10$ to strengthen the L2 penalty in the loss function. Furthermore, minimum performance thresholds for accuracy and F1-score were enforced based on the performance observed for each group when applying a static threshold of 0.5 and a ROC-based threshold. As stricter fairness constraints are imposed, these minimum thresholds are gradually relaxed.
H.3 Hyperparameter settings for the FairOPT
The application of the FairOPT is configured with a set of hyperparameters designed to ensure a balance between fairness and performance during the threshold optimization process. The learning rate is set to $10^{-3}$ to enable controlled and stable updates during optimization. The maximum number of iterations is specified as $10^{5}$ to provide sufficient time for the optimization algorithm to converge. An acceptable disparity of 0.2 is defined to regulate subgroup fairness, ensuring that performance differences across subgroups remain within an acceptable range through the relaxed fairness. Minimum thresholds for accuracy (0.25) and F1-score (0.25) are enforced to maintain baseline model performance and stability, particularly in cases of imbalanced datasets. A tolerance level of $1/2*10^{-4}$ is introduced to impose convergence criteria, ensuring the optimization process halts only when the changes in the objective function are negligible. Furthermore, a penalty parameter of 20 is applied to enforce performance constraints by penalizing significant performance disparities across subgroups. For fairness criteria, Demographic Parity (DP) was employed to test whether the generated thresholds meets the expected fairness. Also, this experiment used early stopping method for effective processing. The early stopping check through the change of the ACC, F1, and maximum DP discrepancy of each group, and terminates if all of them shows minimal change based on predefined patience.